aramondi:

Calculate the number of grams of Fe2O3 needed to react with 19.0 g C.

Vocaloid:

Does it give you an equation? Because a redox reaction can occur between C and Fe2O3 to make CO, or CO2 (although I believe CO2 would be favored)

surjithayer:

at. mass of Fe=56 at. mass of C=12 at. mass of O=16 \[2Fe_{2}O _{3}+3C=4 Fe+3CO_{2}\] 2(56*2+3*16)+3*12=... 2(112+48)+36=... 2*160+36=... 320+36=... 36 gms of C react with =320 gms. of[Fe_{2}O_{3}\] 19 gms of C react with =320/36 *19 =?....